AO Com Work: 26 October 2010 UT

Observer: DMiller, SEsposito, MXompero, LBusoni, APuglisi, FQuiros
Telescope Operator: DHuerta (LBTO)
Telescope Support: JUrban (LBTO)
Instrument Support: TShih (LBTO)

Summary

Preparations

Details

BP4 ACTIVE closed dome tests

02:20 David shut down all GUI's, Subsystems and network server. He the switched the build to BP4 via the lbtswitcher. He also logged in to tcs5 and used lbtswitcher to change to build BP4

02:30 David restarted all networks server (from TCS GUI), but we noticed MCS was still running and there were several GUI's still running. So he shutdown all network servers, we logged into all machine, kill MCS and all GUI's.

02:40 David started all network server

02:47 Doug set GCS into simulation mode by setting noAGW and noAZcam to true in lbt.conf

02:50 David stopped and started GCSR

02:55 preset, M5_0881, gs=0, MODE=TRACK Success

02:57 preset, M5_0881, gs=0, MODE=ACTIVE Failed. GCSR is not in simulation mode and took an image with AZcam

03:00 Checked lbt.conf. I had set GCSL to simulation. Now set GCSR to simulation. David restarted GCSR

03:02 preset, M5_0881, gs=0, MODE=ACTIVE Success. Find a star in acquisition image and guide and wfsc

BP4 ADAPTIVEACE closed dome tests

03:02 preset, M5_0881, gs=0, MODE=ADAPTIVEACE Failed. ccd47 in wrong state

03:02 preset, M5_0881, gs=0, MODE=ADAPTVEACE Success, including presetAO

03:10 Alfio reset some stuff in AO supervisor to allow closed loop with flowerpot fiber (W unit test source) Retroreflector is not mounted

03:15 preset, M5_0881, gs=0, MODE=ADAPTVEACE Success

03:15 Michele called

03:27 irc_runao, false Failed. IDL file error. IDL did not like Alfio's 0 gain filename

03:29 Alfio changed the name

03:30 irc_runao, false Failed. IDL file error. IDL still did not like Alfio's 0 gain filename

03:32 Alfio changed the name again

03:33 irc_runao, false Failed. IDL file error. IDL still did not like Alfio's 0 gain filename. Actually need to send a new preset so names of new files are read.

03:34 preset, M5_0881, gs=0, MODE=ADAPTVEACE Success

03:36 irc_runao, false Success AO closed loop on the test source. The status boxes on the IIF GUI are all correct

03:39 irc_pauseao Success. The AO system is now paused. The IIF GUI shows AO as Open(Red) This should be Paused(Yellow)...maybe

03:39 irc_resume Success. Loop closed and IIF status is correct (Closed).

03:41 iif_offset, 0.5, 0.0 Success. AO loop is paused, telescope slewed, AO loop is resumed. Is the GSC probe position also changed to reflect this offset? I could not tell on the GUI. Will try again later while watching the AGW probe position on the agw-control computer.

03:44 iif_preset, M5_0881, gs=0, MODE=ACTIVE Success This was to open the AO loop

Alfio says there is an offset without the pause and resume from IIF. Is this implemented?

Supper

04:49 iif_preset, M5_0881, gs=2, MODE=ACTIVE Success During closed dome test last week, off-axis guide stars caused a big slew and failure. May have been too far off-axis (outside the guide probe range, but GCS simulator did not care)

04:52 iif_preset, M5_0881, gs=2, MODE=ACTIVE, POSANGLE=180 Success. The guide probe can not move this position, but still successfull

05:10 Test WFSC "Power Off" button on AOS GUI. Successful

05:11 Test ADSEC "Power Off" button on AOS GUI. Successful

05:12 Test WFSC "Power On" button on AOS GUI. Successful. Took about 5 minutes

05:20 Test ADSEC "Power On" button on AOS GUI. Successful. Took about 1.5 minutes. Shell is against reference body.

05:23 Test ADSEC "Set" button on AOS GUI. Successful. Took about 30 seconds

05:30 David reactivated AO5. Took about 10 minutes from start of switch to ready for a preset. Quite Good!

05:40 Ready for a preset. Still closed due to high winds: 20 m/s

Open

Humidity 25%, wind 16m/s

08:45 Open the chamber

08:51 Slew to a BS star and David acquires a star and set IE=-36 CA=37

09:00 pointto BS* in ACTIVE mode. Collimate to 500 nm rms

09:10 iif_preset, AO_04p00m08_2, gs=2, MODE=ACTIVE, /YAXIS

09:12 pointto M5_0162 ACQUIRE find star with IE=-75 CA=46

09:20 iif_preset, AO_04p00m08_2, gs=2, MODE=ACTIVE, /YAXIS

09:24 Collimated and AO team takes over.

A bunch of AO data on this star

12:15 iif_preset, AO_08p00m08_0, gs=4, MODE=ACTIVE Failed with no guide star

We then went an M5 star and set IE and C

iif_preset, AO_08p00m08_0, gs=4, MODE=ACTIVE Failed with no guide star again. Clear Active Optics and thin shell corrections on Secondary hexapod. preset again, but still nothing

We then went another M5 star and set IE and C

12: 45 iif_preset, AO_08p00m08_0, gs=4, MODE=ACTIVE Success

Collimated for a few minutes and then AO Team took over

13:00 Closed

==> Note, presetting to a new star should not take 30 minute. We need to clear Active Optics and Thin Shell corrections from M2 hexapod before slewing to a new star

==> Collimation is taking 10 iterations because the there is not lookup table being applied to the thin shell due to the gravitational induced astigmatism. We really need a lookup table. Can on be applied to the primary? We need Z6.

-- DougMiller - 26 Oct 2010
Topic revision: r5 - 27 Oct 2010, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback